OpenStudy (anonymous):

what is x to the power of 2 (x) x to the power of 1

zepdrix (zepdrix):

Hey Karleigh! :) Don't use x for multiply (that's so confusing!!). So I'm assuming the problem is: `x to the power of 2` times `x to the power of 1` x^2 * x^1 \(\Large\rm x^2\cdot x^1\) When we multiply stuff with the same base, we `add` the exponents.\[\Large\rm x^2\cdot x^1=x^{2+1}\]